I can't seem to figure out how to work this problem:
Find the volume of a cylinder, do not approximate pi.
r = (y-3)cm
h = 7ycm
I've tried:
y^2-6*7= 7y^2-42
and
ended up with y^2-3y-9

Answer by checkley77(12569)

r = (y-3)cm
h = 7ycm
V=pir^2h
V=22/7*(y-3)^2*7y
V=22/7(y^2-6y+9)7y
V=22/7*7y(y^2-6y+9)
V=22y(y^2-6y+9) ans.